OpenStudy (anonymous):

rationalize sqrt((sqrt3-1)/(2sqrt3))

OpenStudy (anonymous):

\[\frac{\sqrt{\sqrt{3}-1}}{2\sqrt{3}}\]?

OpenStudy (anonymous):

yes